OpenStudy (anonymous):

A car is traveling at a speed of 90 miles per hour. What is the cars speed in kilometers per hour? How many kilometers will the car travel in 3 hours? In your computations assume that 1 mile is equal to 1.6 kilometers. Do not round your answer.

OpenStudy (johnweldon1993):

So if we are given that 1mi is equal to 1.6km We have \[\large 90\frac{\cancel{mi}}{h}\times (\frac{1.6km}{1\cancel{mi}}) = ?\frac{km}{h}\]

OpenStudy (anonymous):

Thank you. I understand it now.. 144 km/h 3 hours 432 km

OpenStudy (johnweldon1993):

Good job :)